ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

9. Ordinary Differential Equations and Dynamical Systems

1991 MS Classification: 76 Makinde, Oluwole Daniel Mathematics Department, University of Zimbabwe, Harare, Zimbabwe Numerical Study of Bifurcation for Steady Flow in a Diverging Tube The steady flow in a diverging tube of small aspect ratio is investigated. We obtained the governing non-linear differential equation involving small parameter Re (the flow Reynolds number). Using perturbation technique and computing aid (MAPLE), the solution for the various flow characteristics are obtained and the bifurcation study is performed (Makinde, 1997). The study reveals accurately the Reynolds number at which the separation and reversal in the flow field occur. Reference: O. D. Makinde, Steady flow in a linearly diverging asymmetrical channel, CAMES 4 (2) (1997), 157-167.